OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 4.12(x – 7.89) = 18.12 A. 12.4 B. 12.5 C. 12.2 D. 12.6

OpenStudy (anonymous):

@Death_The_Kid can you help me out with this one please

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Ok I'm going to have you work it out with me this time because I explained it all the last time, ok?

OpenStudy (anonymous):

1. identify. What are we working with, and what are the possible answers?

OpenStudy (anonymous):

2. we will substitute the answer choices in for x. (example: 3(7-x)=2 and instead of x, we would substitute an answer choice. 3(7-4)=2

OpenStudy (anonymous):

3. round. we will round the numbers (not the substituted answer choices) either up or down. (1-4=down 5+=up. don't worry about the extra numbers behind it.)

OpenStudy (anonymous):

if you were doing it with me, your paper should look something like this 

OpenStudy (anonymous):

then you plug it into the calculator to find the correct answer. Now, you tell ME what the answer is!

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

not quite try again!

OpenStudy (anonymous):

Oh I think it's 12.5

OpenStudy (anonymous):

Yes! Good work!

OpenStudy (anonymous):

Yay, thank you for helping me out!

OpenStudy (anonymous):

I encourage you take notes on the things you learn on openstudy. I can't make you, but remember that during PASS testing, you won't be able to just ask someone do the answers.

OpenStudy (anonymous):

Youre welcome!

OpenStudy (anonymous):

alright i will!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

instead of using a regular notebook, i use microsoft word or textedit, and copy/paste my notes into there from openstudy or anywhere else and save it to my desktop in a school folder.